# TUN: Detecting Significant Points in Persistence Diagrams with Deep Learning

### v1

@@ I Introduction
Topological Data Analysis (TDA), particularly through the framework of persistent homology, has emerged as a powerful mathematical approach for extracting topological information from complex data. By constructing multi-scale representations of data, persistent homology robustly quantifies essential topological features such as connected components, loops, and voids, summarizing them in concise descriptors known as persistence diagrams (PDs) [edelsbrunner2010computational]. A point in a PD represents the birth time and death time of a topological feature, and a brief introduction of the mathematical foundations of TDA, particularly filtration, persistent homology, and persistence diagram, is presented in Appendix -A.
A central challenge within TDA is achieving accurate topology understanding, particularly for widely used point cloud representations. Point clouds have become a fundamental representation of three-dimensional objects and environments, with widespread applications across scientific and engineering disciplines, including robotics [zhang2014loam], medical imaging [singh2023topological], and computational biology [xia2014persistent]. Topology understanding is crucial for processing geometric models represented by point clouds, as the topology can have a decisive impact on the model’s functionality. When a point cloud is sampled from a geometric shape, its PD holds the potential to reveal intrinsic topological invariants (for example, the genus, the Betti number, and the loop or void features), providing insights into the fundamental structure of the original object. Accurate topology understanding can help us perform more reliable analysis or performance prediction of the model.
@@
However, the promise of automated topological discovery from persistence diagrams remains largely unfulfilled. The interpretation of PDs constitutes a challenging bottleneck in the TDA pipeline. A PD is a multiset of birth-death pairs, and the core problem involves distinguishing points corresponding to genuine, significant topological features from those representing topological noise or minor geometric details. Traditional heuristics often assume that points with high persistence (long lifespan) are significant [fasy2014confidence], while those near the diagonal represent noise. Yet, this simple thresholding approach proves unsatisfactory in practice [bubenik2020persistent], as the concept of “significance” is inherently context-dependent. Accurate interpretation requires comprehensive assessment, considering not only persistence but also absolute birth and death times, relationships to the global distribution of other points in the diagram, and crucially, the geometric properties of the original point cloud. This complex analysis has traditionally relied on manual expertise, demanding a deep understanding of algebraic topology and rendering the process subjective, time-consuming, and difficult to scale.
The primary goal of this study is to interpret one-dimensional persistence diagrams (1-PDs) computed by alpha filtration [edelsbrunner2010computational] so as to recover the genuine 1-d loop features of the underlying geometric shapes from which the point clouds are sampled. These loop features can be essential for downstream tasks such as property prediction and control. Accordingly, a point in a 1-PD is deemed significant only if it corresponds to such a true 1-d loop of the original model. Because the topology of these shapes can be intricate, the decision must exploit both the internal structure of the persistence diagram and the geometric information carried by the point cloud itself. Classic persistence-threshold methods [fasy2014confidence, he2023robust], which usually ignore the link between diagram features and the topological invariants of the shape, are insufficient for this purpose. We give a brief introduction to some classic methods for topology understanding of PD in Appendix -A.
@@
To address these challenges, we present Topology Understanding Net (TUN), a deep learning framework that formulates topology understanding of PD as per‑point classification: each diagram point is labeled as significant or noise. TUN uses a multi‑modal architecture that fuses topological descriptors from persistence diagrams with geometric context from the raw point cloud, enabling the model to resolve ambiguities that PD‑only heuristics cannot by learning how geometry and topology jointly determine feature significance. In this study, we focus on one‑dimensional persistence diagrams (1‑PDs), where identifying significant points is essential for understanding loop patterns in curves and surfaces. In our setting, 1‑PDs are computed from the alpha filtration [edelsbrunner2010computational]; significant points correspond to generators of the first homology group, and their count equals the first Betti number. The formulation can naturally extend to higher‑dimensional diagrams. However, a major obstacle for this task has been the absence of datasets for supervised learning. We therefore construct the first large‑scale dataset for topology understanding, providing ground‑truth labels for significant points in computed 1‑PDs across diverse categories of point clouds. Experiments show that TUN consistently outperforms traditional methods and attains near‑perfect accuracy, offering a practical tool that bridges abstract topological representations with concrete geometric data for large‑scale analysis.
In summary, our main contributions are as follows:
  • We construct and release the first large-scale, manually labeled dataset specifically designed for supervised learning of topological feature significance in 1-PDs.
  • We propose Topology Understanding Net (TUN), the first deep learning framework for automated topology understanding of persistence diagrams, which uniquely leverages both topological and geometric data through a multi-modal architecture.
  • Experimental results demonstrate that TUN achieves state-of-the-art performance, significantly outperforming traditional methods across four diverse categories of datasets.
The remainder of this paper is organized as follows. Section II reviews prior work on topological data analysis and classical approaches to interpreting persistence diagrams. Section III details the proposed TUN model. Section IV evaluates TUN on four benchmark datasets and provides a comparative analysis. Section V concludes this paper.
@@ II Related Work
II-A Topological Data Analysis
Topological Data Analysis (TDA) [edelsbrunner2008persistent, edelsbrunner2010computational, zomorodian2004computing, wasserman2018topological] offers mathematically grounded tools to study the shape of data in a way that is robust to noise and invariant to many sampling artifacts. The central construct is persistent homology, which tracks the birth and death of topological features across a filtration and summarizes them in persistence diagrams (PDs). Foundational texts and surveys have established theoretical properties (such as the computational methods of PDs for different kinds of filtrations [edelsbrunner2010computational, maria2014gudhi] and the stability of PDs under perturbations [cohen2005stability]) and have documented successful applications across disciplines, such as molecular science [xia2014persistent, bramer2020atom, nguyen2022topological, xia2023persistent], materials science [xia2015persistent, hiraoka2016hierarchical, lee2017quantifying, obayashi2022persistent, suzuki2023persistent], cognitive science [liu2021persistent, chen2025gestalt], geometric design and processing [bruel2020topology, dong2022topology, yan2023reasonable, he2023robust, he20243d, chen2025robust, gao2025persistent], and medical image processing [brito2025persistent, singh2023topological].
In particular, TDA has been integrated with machine learning frameworks to enhance feature extraction and classification tasks [pun2022persistent, barnes2021comparative]. For example, PDs can be integrated into a deep learning framework as a topological prior. PersLay [carriere2020perslay] embeds PDs into a vector space using learnable functions, enabling them to be processed by standard neural network layers. Persformer [reinauer2021persformer] embeds persistence diagrams into deep learning by treating each diagram as an unordered set of points and utilizing a Transformer to learn hidden features without hand-crafted vectorization. Moreover, many studies integrate PDs into loss functions to give more topology-aware training results [wong2021persistent, hu2025topogen, jignasu2024sdfconnect]
@@ II-B Topology understanding of PD
Understanding the significance of points in PDs has been explored using statistical and geometric approaches. A common heuristic ranks points by persistence and declares the top $k$ (or the single most persistent point) as significant, discarding the remainder as noise [bruel2020topology, dong2022topology]. A more principled statistical alternative constructs a confidence set around the diagonal of a PD via bootstrap resampling to separate signal from noise [fasy2014confidence]. This procedure depends on the underlying point-cloud density and requires user-selected parameters, limiting automation. Clustering-based strategies (e.g., $k$-means with $k=2$) similarly partition PD points into significant and noise classes [he2023robust, chen2025robust], but for cases that a PD contains only significant points or only noise points, this method is not feasible. In summary, these families of methods ultimately rely on persistence as the primary importance measure, which is intrinsically limited [bubenik2020persistent]: persistence may omit the geometric context that generated the diagram. Low-persistence features may be genuinely meaningful when small-scale features are present, whereas high-persistence features may be artifacts of outliers.
Figure 3: The entire network architecture of TUN.
III Methods
In this section, we introduce our novel deep learning framework, the TUN, designed for the automated classification of significance in persistence diagrams (PDs). The model leverages a multi-modal architecture that jointly processes persistence diagrams and their corresponding raw point cloud data to deliver robust and accurate predictions. Our TUN network consists of four main components: a Persistence Diagram Encoder, a Point Cloud Encoder, a Multi-Modal Fusion module, and a final Significance Classifier. The overall architecture is depicted in Figure 3. More details of implementation and hyperparameters of our TUN model are available in Appendix -B.
III-A Input Feature Extraction and Processing
Persistence Diagram Features
Point Cloud Data
Auxiliary Features
III-B Model Architecture
Persistence Diagram Encoder
We encode each persistence diagram as a set of enhanced per-point descriptors and aggregate them using attention [vaswani2017attention], finally obtaining $g_{\text{pd}}$. Given raw pairs $(b_{i},d_{i})$, we construct a 4D point feature $v_{i}=[b_{i},d_{i},p_{i},r_{i}]$ with $p_{i}=d_{i}-b_{i}$ and $r_{i}=\log\tfrac{d_{i}}{b_{i}}$. Numerical stabilization is applied to avoid overflow and degeneracy. Each $v_{i}$ is processed by a point-wise MLP with standard nonlinearities and normalization, followed by a multi-head self-attention layer to model relationships among points. This yields contextualized per-point features $F_{i}$, from which a global diagram feature is obtained via pooling.
Point Cloud Encoder
To incorporate geometric context, we use a point cloud encoder operating directly on raw point clouds, inspired by [qi2017pointnet]. The point branch employs a stack of $1\times 1$ Conv1d layers, followed by normalization and ReLU, to generate features for each point. Subsequently, a compact global branch refines these representations, which are then aggregated into a global geometric descriptor $g_{\text{pc}}$ through a max-pooling operation.
Multi-Modal Fusion
Significance Classifier
III-C Loss Function
We adopt a weighted Focal Loss to address the severe class imbalance between rare significant points and abundant noise points. For each PD point $i$, let $\mathbf{s}_{i}\in\mathbb{R}^{2}$ be the two-class logits, $\mathbf{p}_{i}=\operatorname{softmax}(\mathbf{s}_{i})$ the class probabilities with components $\mathbf{p}_{i}[0],\mathbf{p}_{i}[1]\in[0,1]$, and $y_{i}\in\{0,1\}$ the ground-truth label (1: significant, 0: noise). Denote by $p_{t,i}=\mathbf{p}_{i}[y_{i}]$ the probability assigned to the true class. The per-point focal term is
$\ell_{i}=w_{y_{i}}\,\alpha\,\big(1-p_{t,i}\big)^{\gamma}\,\big(-\log p_{t,i}\big),$
where $w_{y_{i}}\in\mathbb{R}_{+}$ is the class weight of label $y_{i}$ (typically $w_{1}>w_{0}$ to upweight the positive class), $\alpha>0$ is a global scaling coefficient, and $\gamma\geq 0$ is the focusing parameter that down-weights easy examples (large $p_{t,i}$) and emphasizes hard ones. Since PDs are padded to a fixed length, we compute the loss only over valid points via a mask $M_{i}\in\{0,1\}$ and normalize by the number of valid points:
$\mathcal{L}=\frac{1}{\sum_{i}M_{i}}\sum_{i}M_{i}\,\ell_{i}.$
IV Results
TABLE I: Performance comparison of our model (TUN) against baselines and an ablation variant. “TUN” is our full model, “2-means” is a clustering baseline, and “CS” is a confidence set method baseline.
Planar shapes (250 samples) CAD models (500 samples) TPMS (500 samples) Zeolite (250 samples)
Model F1 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec
2-means 0.6455 0.9483 0.5213 0.8474 0.4478 0.9292 0.3153 0.7725 0.4765 0.8231 0.3498 0.7474 0.5866 0.7501 0.4278 0.9327
CS(0.5) 0.2505 0.9521 0.9615 0.1440 0.1239 0.9652 0.9762 0.0662 nan 0.8923 nan 0.0000 nan 0.8099 nan 0.0000
CS(0.9) 0.1825 0.9498 0.9655 0.1008 0.1077 0.9649 0.9725 0.0570 nan 0.8923 nan 0.0000 nan 0.8099 nan 0.0000
TUN 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 0.9997 0.9999 0.9994 1.0000 1.0000 1.0000 1.0000 1.0000
TABLE II: Detailed performance metrics (F1-score, Accuracy, Precision, Recall) for feature ablations.
Planar shapes CAD models TPMS Zeolite
Model F1 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec
Ablation 1 0.9837 0.9982 0.9679 1.0000 0.9947 0.9996 0.9900 0.9995 0.9951 0.9989 0.9908 0.9994 0.9911 0.9966 0.9824 1.0000
Ablation 2 0.9904 0.9989 1.0000 0.9809 0.9968 0.9998 1.0000 0.9936 0.9982 0.9996 0.9994 0.9969 0.9993 0.9997 1.0000 0.9985
Ablation 3 0.9968 0.9996 1.0000 0.9936 0.9995 0.9999 1.0000 0.9989 0.9984 0.9997 0.9994 0.9974 0.9998 0.9999 1.0000 0.9996
Ablation 4 0.9964 0.9996 1.0000 0.9929 0.9992 0.9999 0.9984 1.0000 0.9995 0.9999 0.9991 1.0000 1.0000 1.0000 1.0000 1.0000
Ablation 5 0.9978 0.9998 1.0000 0.9957 0.9997 0.9999 1.0000 0.9995 0.9985 0.9997 0.9994 0.9976 0.9998 0.9999 1.0000 0.9996
Ablation 6 0.9975 0.9997 1.0000 0.9950 0.9995 0.9999 1.0000 0.9989 0.9986 0.9997 0.9994 0.9978 0.9999 0.9999 1.0000 0.9998
TUN 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 0.9997 0.9999 0.9994 1.0000 1.0000 1.0000 1.0000 1.0000
IV-A Datasets
We evaluate on four categories of point-cloud data: (i) planar shapes; (ii) CAD models drawn from the ABC dataset [koch2019abc] and Thingi10k dataset [zhou2016thingi10k]; (iii) triply periodic minimal surface (TPMS) structures [wang2024simultaneous]; and (iv) zeolite crystal frameworks [baerlocher2007atlas]. For each instance, we construct a point cloud by sufficiently sampling the underlying geometric model and, when specified, perturbing points with additive noise at different levels. Some examples are illustrated in Figure 7 in Appendix -C. We then compute its one-dimensional persistence diagram. Based on the identification rule of significant points discussed in previous sections, significance labels are assigned by category: for planar shapes, we annotate significant points via direct counting-based inspection of the diagram; for 3D models whose sampled surfaces form 2D manifolds, we compute the genus $g$, derive the one-dimensional Betti number $\beta_{1}$ ($\beta_{1}=2g$ for closed orientable surfaces), and label the diagram’s significant birth-death pairs in accordance with $\beta_{1}$; for surfaces with boundary or multiple connected components, we compute Betti numbers by evaluating homology group ranks based on homology theory [hatcher2002algebraic] and use these to guide the labeling. Owing to sufficient sampling, significant points typically emerge at early filtration times, exhibit large persistence, and their count matches the first Betti number $\beta_{1}$ of the underlying geometric model.
The dataset consists of 3,000 training samples and 500 validation samples (totaling 3,500) and is distributed among four classes; the test set contains 1,500 samples: 250 planar shapes, 500 CAD models, 500 TPMS, and 250 zeolites. Each point cloud is sampled from a shape whose first Betti number is fewer than 100, guaranteeing fewer than 100 significant points in the corresponding 1-PD. Point counts per point cloud range from hundreds to 100,000, ensuring diversity in density and noise level. Each PD sample is either truncated or padded to consist of 100 points, and each point cloud is subsampled or padded to contain 50,000 points.
IV-B Evaluation Metrics
We evaluate the performance of our model using standard classification metrics, including accuracy (Acc), precision (Pre), recall (Rec), and F1-score. The required TP, TN, FP, FN metrics for computing these metrics are given in Appendix -D. The best model is selected based on the lowest loss achieved on a held-out validation set, and the main evaluation metric for testing is F1-score. This choice is motivated by the severe class imbalance in our data: for each PD, we retain the top-100 points with the largest persistence, among which the number of positive labels (significant points) is typically much smaller than that of negative labels (noise points). F1-score, as the harmonic mean of precision and recall, provides a balanced measure that is especially sensitive to the model’s ability to correctly identify the minority positive class, making it the most informative metric for this imbalanced classification task.
IV-C Results and Analysis
We present the quantitative results of our TUN in Table I. The evaluation is conducted on four distinct datasets, comparing with classic methods: the 2-means clustering method and the Confidence Set (CS) method. The 2-means clustering method uses clustering to classify the points in 1-PD into two classes, i.e., the classes of significant points and noise points. And the confidence region method determines a threshold and regards points whose persistence is larger than this threshold as significant. In our experiments, we used GUDHI  [^1] https://github.com/GUDHI/TDA-tutorial/blob/master/Tuto-GUDHI-ConfRegions-PersDiag-datapoints.ipynb[maria2014gudhi] tools to calculate the confidence set for each 1-PD. We tested with confidence levels of 0.5 and 0.9 (denoted as CS(0.5) and CS(0.9), respectively), while leaving other parameters at their default values. Additionally, we also compare our TUN full model against ablation variants that rely solely on persistence diagram features or exclude the auxiliary features.
Analysis and Comparison.
Our full model, TUN, demonstrates exceptional and state-of-the-art performance across all four datasets. As shown in Table I, it achieves a perfect F1-score, accuracy, precision, and recall of 1.0000 on the Planar Shapes, CAD Models, and Zeolite datasets, indicating flawless classification. On the more intricate TPMS dataset, it attains a near-perfect F1-score of 0.9997. These results establish the superiority and robustness of our deep learning-based, multi-modal approach, confirming that TUN can consistently and accurately identify true topological features in point clouds of varying complexity. Additionally, when comparing with traditional methods, the performance of these methods highlights the difficulty of the task and the limitations of heuristic-based approaches. The 2-means clustering algorithm [he2023robust] yields poor F1-scores, dropping as low as 0.4436 on CAD models. Its low precision across all datasets indicates that it incorrectly classifies many noise points as significant, failing to provide a reliable interpretation. The Confidence Set method [fasy2014confidence] performs even more poorly in terms of F1-score. While it can achieve high precision on simple data by being extremely conservative, its recall is exceptionally low (e.g., 0.2505 for CS(0.5) on Planar Shapes). On the more complex TPMS and Zeolite datasets, its recall drops to zero, causing the F1-score to be undefined (nan). This demonstrates that simple persistence thresholding is brittle and often fails on challenging data.
Ablation Studies.
Moreover, we conduct ablation studies to rigorously evaluate the contribution of each component in our multi-modal framework. We introduce several ablation variants: (1) Ablation 1: Model only has the persistence diagram encoder, without the point cloud encoder and any auxiliary features. (2) Ablation 2: TUN model without all auxiliary features, only has persistence diagram encoder and point cloud encoder. (3) Ablation 3: TUN model without the 5 persistence diagram statistics. (4) Ablation 4: TUN model without the 3 point cloud statistics. (5) Ablation 5: TUN model without the 3 noise and uniformity estimation features. (6) Ablation 6: TUN model without the 3 bounding box features. Based on Table II, the ablations reveal how each component contributes. “Ablation 1” (PD encoder only) already outperforms traditional baselines, with F1 of 0.9837 / 0.9947 / 0.9951 / 0.9911 on Planar / CAD / TPMS / Zeolite. However, its precision is comparatively low, indicating it generates some false positives, while its recall is near 1.0000 across all datasets, showing strong sensitivity to significant features. “Ablation 2” adds point-cloud geometry and significantly improves precision (reaching 1.0000 on Planar / CAD / Zeolite and 0.9994 on TPMS), effectively filtering out noise. This comes with a slight trade-off in recall, suggesting the model becomes slightly too conservative and misses some subtle features. “Ablation 3” removes global PD statistics, which results in a decrease in recall while maintaining high precision. This implies that global diagram summaries help the model recover significant points that might otherwise be ignored. “Ablation 4” removes point-cloud statistics, similarly leading to a drop in recall with robust precision. Geometric aggregates prove important for maintaining the model’s sensitivity to genuine topological features. “Ablation 5” discards noise and uniformity estimation features, causing recall to decline while precision remains almost unaffected. These features assist the model in identifying significant structures without increasing the false positive rate. “Ablation 6” removes bounding-box features and exhibits a similar trend, with recall lowering while keeping precision high. Spatial extent cues help regularize decisions, preventing the omission of valid features. Despite the strong performance of ablation variants, the full TUN model consistently optimizes both precision and recall, achieving near-perfect F1 and accuracy across all datasets.
Figure 4: Illustrating a failure of TUN model. In this 1-PD of TPMS test sample (ID:000491), three points corresponding to small-scale loops in the original model have small persistence (marked in red in Ground Truth, while blue in Prediction), which makes them hard to distinguish from noise points.
IV-D Limitations
In this study, we restrict the number of significant points in 1-PD to fewer than 100 and the number of point cloud points to fewer than 100,000. This is due to the consideration of computational efficiency based on the assumption that the significant points are no more than 100. If the number of points in the 1-PD and the point cloud is too large, the training process will be time-consuming and memory-intensive. But theoretically, our TUN model allows handling larger datasets with larger point clouds and PD with more points. Additionally, small-scale loops are challenging for TUN to detect reliably. For instance, in the TPMS dataset, our TUN model inaccurately predicted one sample that includes three small-scale loops, and their associated points in the 1-PD exhibit small persistence, resembling noise points (see Figure 4). One possible solution is to implement an extremely dense sampling strategy, but executing dense sampling to ensure the detection of these loops could result in high computational costs. Consequently, finding methods to accurately capture such features remains a challenge.
V Conclusion
In future work, we will explore the topology understanding of higher-dimensional PDs, as well as explore alternative inputs instead of point clouds (like images, voxels) to broaden the applicability and improve the interpretability of the model.
-A Mathematical Foundations of TDA

### v2

@@ I Introduction
Topological Data Analysis (TDA), particularly through the framework of persistent homology, has emerged as a powerful mathematical approach for extracting topological information from complex data. By constructing multi-scale representations of data, persistent homology robustly quantifies essential topological features such as connected components, loops, and voids, summarizing them in concise descriptors known as persistence diagrams (PDs) [13]. A point in a PD represents the birth time and death time of a topological feature.
A central challenge within TDA is achieving accurate topology understanding, particularly for widely used point cloud representations. Point clouds have become a fundamental representation of three-dimensional objects and environments, with widespread applications across scientific and engineering disciplines, including robotics [44], medical imaging [34], and computational biology [42]. Topology understanding is crucial for processing geometric models represented by point clouds, as the topology can have a decisive impact on the model’s functionality. When a point cloud is sampled from a geometric shape, its PD holds the potential to reveal intrinsic topological invariants (for example, the genus, the Betti number, and the loop or void features), providing insights into the fundamental structure of the original object. Accurate topology understanding can help us perform more reliable analysis or performance prediction of the model.
@@
However, the promise of automated topological discovery from persistence diagrams remains largely unfulfilled. The interpretation of PDs constitutes a challenging bottleneck in the TDA pipeline. A PD is a multiset of birth-death pairs, and the core problem involves distinguishing points corresponding to genuine, significant topological features from those representing topological noise or minor geometric details. Traditional heuristics often assume that points with high persistence (long lifespan) are significant [14], while those near the diagonal represent noise. Yet, this simple thresholding approach proves unsatisfactory in practice [6], as the concept of “significance” is inherently context-dependent. Accurate interpretation requires comprehensive assessment, considering not only persistence but also absolute birth and death times, relationships to the global distribution of other points in the diagram, and crucially, the geometric properties of the original point cloud. This complex analysis has traditionally relied on manual expertise, demanding a deep understanding of algebraic topology and rendering the process subjective, time-consuming, and difficult to scale.
The primary goal of this study is to interpret one-dimensional persistence diagrams (1-PDs) computed by alpha filtration [13] so as to recover the genuine 1-d loop features of the underlying geometric shapes from which the point clouds are sampled. These loop features can be essential for downstream tasks such as property prediction and control. Accordingly, a point in a 1-PD is deemed significant only if it corresponds to such a true 1-d loop of the original model. Because the topology of these shapes can be intricate, the decision must exploit both the internal structure of the persistence diagram and the geometric information carried by the point cloud itself. Classic persistence-threshold methods [14, 17], which usually ignore the link between diagram features and the topological invariants of the shape, are insufficient for this purpose.
@@
To address these challenges, we present Topology Understanding Net (TUN), a deep learning framework that formulates topology understanding of PD as per‑point classification: each diagram point is labeled as significant or noise. TUN uses a multi‑modal architecture that fuses topological descriptors from persistence diagrams with geometric context from the raw point cloud, enabling the model to resolve ambiguities that PD‑only heuristics cannot by learning how geometry and topology jointly determine feature significance. In this study, we focus on one‑dimensional persistence diagrams (1‑PDs), where identifying significant points is essential for understanding loop patterns in curves and surfaces. In our setting, 1‑PDs are computed from the alpha filtration [13]; significant points correspond to generators of the first homology group, and their count equals the first Betti number. The formulation can naturally extend to higher‑dimensional diagrams. However, a major obstacle for this task has been the absence of datasets for supervised learning. We therefore construct the first large‑scale dataset for topology understanding, providing ground‑truth labels for significant points in computed 1‑PDs across diverse categories of point clouds. Experiments show that TUN consistently outperforms traditional methods and attains near‑perfect accuracy, offering a practical tool that bridges abstract topological representations with concrete geometric data for large‑scale analysis.
In summary, our main contributions are: (1) We construct and release the first large-scale, manually labeled dataset specifically designed for supervised learning of topological feature significance in 1-PDs. (2) We propose Topology Understanding Net (TUN), the first deep learning framework for automated topology understanding of persistence diagrams, which uniquely leverages both topological and geometric data through a multi-modal architecture. (3) Experimental results demonstrate that TUN achieves state-of-the-art performance, significantly outperforming traditional methods across four diverse categories of datasets.
@@ II Related Work
II-A Topological Data Analysis (TDA)
Topological Data Analysis  [12, 13, 46, 38] offers mathematically grounded tools to study the shape of data in a way that is robust to noise and invariant to many sampling artifacts. The central construct is persistent homology, which tracks the birth and death of topological features across a filtration and summarizes them in persistence diagrams (PDs). Foundational texts and surveys have established theoretical properties (such as the computational methods of PDs for different kinds of filtrations [13, 27] and the stability of PDs under perturbations [10]) and have documented successful applications across disciplines, such as molecular science [42, 3, 29, 41], materials science [40, 20, 25, 30, 35], cognitive science [26, 8], geometric design and processing [5, 11, 43, 17, 18, 9, 15], and medical image processing [4, 34].
In particular, TDA has been integrated with machine learning frameworks to enhance feature extraction and classification tasks [31, 2]. For example, PDs can be integrated into a deep learning framework as a topological prior. PersLay [7] embeds PDs into a vector space using learnable functions, enabling them to be processed by standard neural network layers. Persformer [33] embeds persistence diagrams into deep learning by treating each diagram as an unordered set of points and utilizing a Transformer to learn hidden features without hand-crafted vectorization. But these frameworks cannot be used to give point-wise significance prediction directly. Moreover, many studies integrate PDs into loss functions to give more topology-aware training results [39, 21, 22]
@@ II-B Topology understanding of PD
Understanding the significance of points in PDs has been explored using statistical and geometric approaches. A common heuristic ranks points by persistence and declares the top $k$ (or the single most persistent point) as significant, discarding the remainder as noise [5, 11]. A more principled statistical alternative constructs a confidence set around the diagonal of a PD via bootstrap resampling to separate signal from noise [14]. This procedure depends on the underlying point-cloud density and requires user-selected parameters, limiting automation. Clustering-based strategies (e.g., $k$-means with $k=2$) similarly partition PD points into significant and noise classes [17, 9], but for cases that a PD contains only significant points or only noise points, this method is not feasible. In summary, these families of methods ultimately rely on persistence as the primary importance measure, which is intrinsically limited [6]: persistence may omit the geometric context that generated the diagram. Low-persistence features may be genuinely meaningful when small-scale features are present, whereas high-persistence features may be artifacts of outliers.
III Preliminaries